# LMGT: Optimizing Exploration-Exploitation Balance in Reinforcement Learning through Language Model Guided Trade-offs*

### v1

@@ Abstract
The uncertainty inherent in the environmental transition model of Reinforcement Learning (RL) necessitates a careful balance between exploration and exploitation to optimize the use of computational resources for accurately estimating an agent’s expected reward. Achieving balance in control systems is particularly challenging in scenarios with sparse rewards. However, given the extensive prior knowledge available for many environments, it is redundant to begin learning from scratch in such settings. To address this, we introduce Language Model Guided Trade-offs (i.e., LMGT), a novel, sample-efficient framework that leverages the comprehensive prior knowledge embedded in Large Language Models (LLMs) and their adeptness at processing non-standard data forms, such as wiki tutorials. LMGT proficiently manages the exploration-exploitation trade-off by employing reward shifts guided by LLMs, which direct agents’ exploration endeavors, thereby improving sample efficiency. We have thoroughly tested LMGT across various RL tasks and deployed it in industrial-grade RL recommendation systems, where it consistently outperforms baseline methods. The results indicate that our framework can significantly reduce the time cost required during the training phase in RL.
I Introduction
Reinforcement Learning (RL) encounters a fundamental challenge in finding a balance between exploration and exploitation [1]. This equilibrium is crucial for ensuring the robustness of RL algorithms when applied in real-world scenarios [2]. Agents operating in such settings often face the exploration-exploitation dilemma due to the unknown and stochastic nature of their environments, making it impossible to deduce the exact environmental model. The interactions between the agent and the environment provide estimates of expected rewards, denoted as , for different actions. However, estimating the mean of rewards using sample means introduces inherent error, which prevents us from being certain that the action with the highest is truly optimal. Consequently, exploration is necessary to reduce the discrepancy between the sample mean and the true mean. However, in situations where resources are limited, it is likely that the action with the highest is indeed the best choice. Accurately estimating the optimal action is a fundamental aspect of RL [3, 4], whereas suboptimal actions require less computational evaluation, reflecting the concept of “exploitation” in RL. The conflict between exploration and exploitation therefore necessitates the selection of actions that are both “close to the current estimate of the best action” while also being “different from it”.
@@ I Introduction
Numerous studies have attempted to address this challenge. Common strategies include -greedy [5], Softmax [5], upper confidence bound [6], thompson sampling [7]. The choice of an appropriate exploration-exploitation strategy depends on the specific application context and problem requirements, as different RL problems may require different strategies. Nevertheless, given the broadening scope of RL applications, manually selecting different exploration-exploitation strategies for each distinct environment is impractical. One major difficulty arises from multimodal and long-tailed data distributions. Some adaptive algorithms adjust the exploration-exploitation balance based on the agent’s experiences [8], but these algorithms still have constraints that can significantly impact model performance and robustness when applied beyond their scope. Additionally, previous exploration-exploitation balance strategies either rely solely on adjusting the ratio based on the data distribution, without utilizing prior knowledge or require the design and adjustment of strategies based on domain expertise and a deep understanding of the task. The latter approach requires substantial manual effort and can potentially reduce learning performance if not properly designed.
To address the limitations of traditional exploration-exploitation strategies, we propose a novel framework called Language Model Guided Trade-offs (i.e., LMGT), that leverages prior knowledge from various sources to guide agents’ inefficient learning with limited resources. Jake et al. [9] discovered that valuable insights can be derived from effective offline demonstration data, enabling agents to align themselves correctly. By capturing the patterns of a sound policy in a model and using this model for intrinsic motivation, the RL agent can effectively map itself to a skillful demonstration-defined subspace. In this subspace, even undirected exploration can significantly enhance the agent’s understanding of the environment if the deviations align with rational pathways. Our LMGT leverages the text comprehension and generation capabilities of Large Language Models (LLMs) to incorporate prior knowledge, thereby enhancing the agent’s environmental understanding and achieving a balance between exploration and exploitation. Leveraging the powerful language processing capabilities of LLMs, our framework obviates the need for highly structured prior knowledge, thereby enabling extensive use of existing human knowledge bases. This distinct advantage sets our approach apart from other methods. Moreover, the text generated by LLMs often reflects the structural patterns of the real world, embedding common-sense knowledge about various aspects of human reasoning and intuition [10, 11]. Compared to some recent methods [12, 13, 14] that directly use LLMs as agents within the RL process, the advantage of our approach lies in the fact that LLMs are only required during the training phase to assist the agent in learning. This fundamentally mitigates the risks associated with LLM hallucinations by confining such risks to the training stage, thereby enhancing the security of the strategy implemented by the agent. Once the training is complete, our agent can be deployed independently without LLMs. In contrast to agents utilizing LLMs kernels, conventional RL agents founded upon multilayer perceptrons or convolutional neural networks exhibit a comparative advantage regarding computational resource utilization [15]. The potential advantages of our architecture become apparent in large-scale application scenarios and latency-sensitive environments. Our interactive process between LLMs and agents involves LLMs processing environmental information and scoring agent behavior to guide exploration and exploitation through reward-shifting mechanisms. Additionally, our method aligns with a key principle: reward shifting is equivalent to modifying the initialization of the Q-function, effectively balancing the exploration and exploitation aspects of RL[16].
@@
We conducted experiments in various settings and environments, demonstrating that our approach effectively utilizes prior knowledge, leading to a reduction in model training costs compared to baseline methods. We also evaluated the performance of different LLMs within our framework, providing a partial assessment of their inferential capabilities. Furthermore, we applied our framework to Google’s industrial-grade recommendation algorithm, SlateQ [17]. Our contributions can be summarized as follows:
We propose a novel framework that leverages LLMs to balance exploration and exploitation within RL. This framework effectively resolves the exploration-exploitation dilemma and provides precise guidance for the agent’s actions. Featuring a distinctive architecture that decouples the LLM from the agent, our approach significantly reduces the risk of LLM hallucinations impacting RL strategies.
@@
We validated our proposed framework in various automatic control environments and RL algorithms. Experimental results demonstrate that our method significantly reduces the training costs of RL models while maintaining both generality and ease of use.
@@
II Methodology
@@ II-B Prompt Design
Table III catalogs a detailed inventory of immediate enhancements utilized in our experimental design. It is important to note that the primary distinction between Zero-shot [18] and Baseline involves Zero-shot’s integration of specific, task-related information into the prompt, which guides the LLM regarding the appropriate information to produce. The Name method could be perceived as perplexing. It involves attributing a name to an LLM in the prompt with the expectation that this modification could enhance its performance. Nonetheless, our experimental results indicate that this technique does not yield any improvements. For additional details on the experiments, please see Section III-C. It is commonplace to deploy multiple prompt enhancements concurrently.
@@
Section III presents an elucidation of the effects of various prompt methods, along with a rationale for our methodological choices.
III Experiment
The experiment is structured into three distinct parts. In the initial phase, we scrutinize the benefits of our proposed framework over conventional approaches for addressing sparse reward challenges. Specifically, we compare LMGT with Return Decomposition for Delayed Rewards (RUDDER) [19], which is a novel RL approach for delayed rewards in finite MDPs. RUDDER’s objective is to neutralize expected future rewards, thereby simplifying Q-value estimations to the average of immediate rewards. Despite RUDDER’s expedited processing in scenarios with delayed rewards compared to traditional RL methods, it fails to incorporate prior knowledge—an area where LMGT particularly excels. Consequently, we anticipate LMGT to facilitate the expedited development of effective behavioral strategies by agents. The second segment of the experiment evaluates our framework’s versatility by applying it across diverse RL algorithms and environments to ascertain its efficacy. This was accomplished by benchmarking across various standard environments provided by Gymnasium [20], an API platform that supports single-agent RL environments such as cartpole, pendulum, mountain-car, mujoco, and atari. Herein, we also examine the enhancement in performance attributable to our framework across various RL algorithms, compared to the baseline algorithms presented in Stable Baselines3[21]—a collection of robust RL algorithm implementations in PyTorch. This phase further includes an assessment of the impact of different prompting techniques on our framework’s performance and an exploratory evaluation of the reasoning capabilities of LLMs within our framework. To ensure that the evaluation conclusions of our framework extend beyond synthetic settings, the final section investigates its practical applications and improvements. Specifically, we explore its integration with Google’s SlateQ [17], a sophisticated recommendation algorithm that employs slate decomposition. This approach effectively manages the complexity of recommending multiple items simultaneously, addressing the challenge of large action spaces found in previous RL recommendation algorithms. This implementation was tested within a simulated environment on RecSim [22], a versatile platform for developing simulation environments for recommender systems (RSs), facilitating sequential user interactions.
III-A Experimental Settings
@@
III-B Comparison Experiments with Traditional Exploration-Exploitation Trade-off Methods
We conducted a comparative study of LMGT and RUDDER using a classic pocket watch repair task as a case study. This task involves repairing and selling a pocket watch, where the decision to repair, contingent on the brand, hinges on cost-benefit analysis given a known selling price versus unknown repair and delivery expenses leading to negative rewards. The challenge lies in delayed rewards, where the profitability of repairing a specific brand becomes apparent only after total costs are established. The objective was to equip the agent with a strategy that consistently achieves a “break-even decision” ratio exceeding . “The number of episodes to learn a qualified strategy” (expressed as Episode) and “the time to learn a qualified strategy” (expressed as Time) served as our evaluation metrics, acknowledging that training time is influenced by hardware performance and hence, primarily offers qualitative insights. In this part of the experiment, to align with the baseline chosen by RUDDER in the example, we applied LMGT to temporal difference (TD). To mitigate random seed effects on outcomes, experiments were conducted using seeds , with results averaged. As evidenced in Table I, LMGT outperformed RUDDER, requiring fewer episodes for strategy acquisition and demonstrating reduced training time. However, the reduced time advantage of LMGT over episodes suggests a potential threshold beyond which the efficiencies derived from LLM guidance might be counterbalanced by computational overheads. This threshold presents a future research direction for us.
| Method | Metric | |
| Episode | Time(sec) | |
| TD | 71823 | 427 |
| MC | 221770 | 530 |
| RUDDER | 2029 | 171 |
| LMGT(ours)+TD | 417 | 114 |

### v2

# Reward Guidance for Reinforcement Learning Tasks Based on Large Language Models: The LMGT Framework

@@ Abstract
The inherent uncertainty in the environmental transition model of Reinforcement Learning (RL) necessitates a delicate balance between exploration and exploitation. This balance is crucial for optimizing computational resources to accurately estimate expected rewards for the agent. In scenarios with sparse rewards, such as robotic control systems, achieving this balance is particularly challenging. However, given that many environments possess extensive prior knowledge, learning from the ground up in such contexts may be redundant. To address this issue, we propose Language Model Guided reward Tuning (LMGT), a novel, sample-efficient framework. LMGT leverages the comprehensive prior knowledge embedded in Large Language Models (LLMs) and their proficiency in processing non-standard data forms, such as wiki tutorials. By utilizing LLM-guided reward shifts, LMGT adeptly balances exploration and exploitation, thereby guiding the agent’s exploratory behavior and enhancing sample efficiency. We have rigorously evaluated LMGT across various RL tasks and evaluated it in the embodied robotic environment Housekeep. Our results demonstrate that LMGT consistently outperforms baseline methods. Furthermore, the findings suggest that our framework can substantially reduce the computational resources required during the RL training phase.
keywords:
Reinforcement Learning , Deep Learning Methods , Learning from Demonstration[1]organization=School of Electronic and Electrical Engineering, Shanghai University of Engineering Science, city=Shanghai, country=China
[2]organization=INFLY TECH (Shanghai) Co., Ltd., city=Shanghai, country=China
1 Introduction
Reinforcement Learning (RL) faces a fundamental challenge in striking an optimal balance between exploration and exploitation [1, 2]. This equilibrium is crucial for ensuring the robustness of RL algorithms in real-world applications [3]. Agents in these environments encounter the exploration-exploitation dilemma due to the inherent unknown and stochastic nature of their surroundings, making it impossible to deduce the exact environmental model.
Through interactions with the environment, agents obtain estimates of expected rewards, denoted as , for various actions. However, estimating rewards using sample statistics inherently introduces error, preventing certainty that the action with the highest is truly optimal. This uncertainty necessitates exploration to reduce the discrepancy between sample means and true means. Conversely, in resource-constrained scenarios, it is often advantageous to select actions with the highest estimated rewards—a practice known as exploitation. The accurate estimation of optimal actions is fundamental to RL [4, 5], whereas suboptimal actions require less computational evaluation. This inherent conflict necessitates selecting actions that balance proximity to the current “best action” estimate while maintaining sufficient diversity [6].
@@ 1 Introduction
Numerous approaches have attempted to address this challenge. Traditional strategies include -greedy [6], Softmax [6], upper confidence bound [7], and Thompson sampling [8]. The selection of an appropriate exploration-exploitation strategy typically depends on the specific application context and problem requirements, as different RL problems demand different approaches. However, with the expanding scope of RL applications, manually selecting distinct strategies for each environment becomes impractical. Particular difficulties arise from multimodal and long-tailed data distributions. While some adaptive algorithms adjust the exploration-exploitation balance based on agent experiences [9], these algorithms still have constraints that can significantly impact model performance and robustness when applied beyond their intended scope. Moreover, previous strategies either rely solely on adjusting ratios based on data distribution without utilizing prior knowledge, or require substantial domain expertise and task understanding, potentially reducing learning performance if improperly designed.
To overcome these limitations, we propose a novel framework called Language Model Guided reward Tuning (i.e., LMGT), which leverages prior knowledge from various sources to guide agents’ learning with limited resources. Based on the work of Bruce et al. [10], who demonstrated that valuable information can be derived from effective offline demonstration data, our approach enables agents to align themselves correctly with desirable behaviors. By capturing patterns of sound policies and using these for intrinsic motivation, RL agents can effectively map themselves to skillful demonstration-defined subspaces where even undirected exploration can significantly enhance environmental understanding if deviations align with rational pathways.
Our LMGT framework harnesses the text comprehension and generation capabilities of Large Language Models (LLMs) to incorporate prior knowledge, thereby enhancing the agent’s environmental understanding and achieving an effective balance between exploration and exploitation. Leveraging the powerful language processing capabilities of LLMs, our framework removes the need for highly structured prior knowledge, thereby allowing broader utilization of human knowledge sources. This distinct advantage sets our approach apart from other methods. Moreover, the text generated by LLMs often reflects structural patterns of the real world, embedding common-sense knowledge about various aspects of human reasoning and intuition [11, 12, 13].
Compared to some recent methods [14, 15, 16, 17, 18, 19] that directly use LLMs as agents within the RL process, the advantage of our approach lies in the fact that LLMs are only required during the training phase to assist the agent in learning. Once training is complete, our agent can be deployed independently without LLMs. In contrast to agents utilizing LLM kernels, conventional RL agents founded upon multilayer perceptrons or convolutional neural networks exhibit a comparative advantage regarding computational resource utilization [20, 21]. These advantages become particularly relevant in large-scale application scenarios and latency-sensitive environments.
Our interaction with LLMs involves LLMs processing environmental information and scoring agent behavior to guide exploration and exploitation through reward-shifting mechanisms. For certain non-text environments, we employ Visual Instruction Tuning [22] to align visual information with the projection matrices of LLMs. The rationale for this approach will be detailed in Section 4.4. Additionally, our method aligns with a key principle: reward shifting is equivalent to modifying the initialization of the Q-function, effectively balancing the exploration and exploitation aspects of RL [23].
@@
We conducted extensive experiments across diverse settings and environments to evaluate our proposed method. The results demonstrate that our approach effectively leverages prior knowledge, significantly reducing the computational resources required for model training compared to baseline methods. Additionally, we assessed the performance of various LLMs within our framework, providing insights into their inferential capabilities in this context. To validate the versatility of our approach, we implemented our framework in Housekeep [24], a simulated robotic environment, and evaluated its performance on complex robotic tasks. Furthermore, we applied our framework to Google’s industrial-grade recommendation algorithm, SlateQ [25].
Our contributions can be summarized as follows:
1.
We propose a novel framework for balancing exploration and exploitation in RL by leveraging LLMs. This framework addresses the exploration-exploitation dilemma and effectively guides the behavior of agents.
2.
@@
We validate our proposed method across various RL environments and algorithms. Our approach significantly reduces the training cost of RL models while maintaining generality and ease of use.
3.
@@
2 Related Work
Our research has significant connections to various preceding studies. Within the realm of reinforcement learning, there has been a comprehensive exploration of strategies to balance exploration and exploitation, a concept that resonates with our methodology. Our approach innovates within this field by harmonizing reinforcement learning principles with the capabilities of LLMs. This harmonization broadens our comprehension and application of balancing strategies, empowering us to tackle decision-making challenges in intricate settings more effectively. Moreover, our study draws inspiration from contemporary research on the evolution of LLMs, which has laid the theoretical groundwork and catalyzed noteworthy advancements in our inquiry.
2.1 Exploration-Exploitation Trade-off in RL
Numerous strategies addressing the RL exploration-exploitation dilemma have been proposed, encompassing various tasks. Established approaches such as the -greedy [6], Softmax [6], Upper Confidence Bound (UCB) [7], and Thompson sampling [8] methods are widely recognized. Furthermore, entropy-based RL frameworks have seen progressive advancements [26, 27], facilitating diverse exploration pathways in Markov Decision Processes (MDPs), rendering them highly effective in continuous action spaces. The emergence of curiosity-driven RL strategies that leverage intrinsic curiosity as a reward signal [28] has spurred extensive research [29]. Notably, these approaches enable agents to interact with their environment and develop new skills. Despite their varying success in numerous contexts, these strategies typically exhibit static features, limiting their adaptability to dynamic settings. An innovative decoupling method was introduced, combining goal-specific and goal-agnostic exploration strategies for improved adaptability [30]. Additionally, deep covering options were investigated to enhance exploration efficiency without reliance on explicit state spaces or unwieldy action sets, courtesy of deep learning models and spectral analysis [31]. Daochen et al. [32] proposed a simplifying exploration method suitable for generative settings, fostering model-independent exploration through a ranking mechanism in environments with sparse rewards. While adept at accommodating changes autonomously, these approaches have restricted applicability due to their inattention to prior knowledge infusion, which could advance the equilibrium between exploration and exploitation. Our framework demonstrates versatility across diverse environments and RL algorithms, incorporating previously untapped prior knowledge to heighten agents’ learning efficacy and minimize resource expenditure during training.
Hao et al. [23] theoretically prove that negative reward shifts aid exploration while positive reward shifts limit exploration. Ingmar et al. [33] introduce a novel reward shaping method that relaxes the strict optimality guarantee of potential-based reward shaping (PB-RS) while maintaining long-term behavioral consistency and enhancing sample efficiency compared to PB-RS. Inspired by these studies, we employed the reward shifts generated by LLMs to balance exploration and exploitation, enabling fine control of the impact on the existing RL process while making the overall learning process more efficient.
2.2 Relevant Research in Large Language Models
LLMs, particularly those based on the Transformer architecture [34], such as OpenAI’s GPT [35, 36], Meta’s Llama [37], Google’s Bard [38], LMSYS Org’s Vicuna [39], and Anthropic’s Claude [40], have widespread applications across various domains. Recent work [41] in the field of LLMs has provided theoretical foundations for our research. [11] demonstrates that the knowledge embedded within LLMs offers valuable inductive biases applicable to both traditional Natural Language Processing (NLP) tasks and non-traditional tasks involving training symbolic reasoning engines. The text generated by LLMs typically reflects structural patterns in the “real world”, indicating that the weights within LLMs encode implicit common-sense knowledge related to various aspects of human reasoning and intuition. Through an analysis of learning representations in the Llama-2 series models across three spatial datasets and three temporal datasets, Wes et al. [12] presents evidence that LLMs form a coherent model of the data-generating process, which suggests that LLMs acquire structured knowledge regarding space and time fundamental dimensions. We posit that the weights embedded within LLMs inherently carry prior knowledge that can be utilized to guide agents’ behavior.
3 Methodology
@@ 3.2 Prompt Design
Table 4 catalogs a detailed inventory of immediate enhancements utilized in our experimental design. It is important to note that the primary distinction between Zero-shot [42] and Baseline involves Zero-shot’s integration of specific, task-related information into the prompt, which guides the LLM regarding the appropriate information to produce. The Name method could be perceived as perplexing. It involves attributing a name to an LLM in the prompt with the aim of improving performance. Nonetheless, our experimental results indicate that this technique does not yield any improvements. For additional details on the experiments, please see Section 4.3. It is commonplace to deploy multiple prompt enhancements concurrently.
@@
Section 4 presents an elucidation of the effects of various prompt methods, along with a rationale for our methodological choices.
4 Experiment
The experiment is structured into three distinct parts. In the initial phase, we scrutinize the benefits of our proposed framework over conventional approaches for addressing sparse reward challenges. Specifically, we compare LMGT with Return Decomposition for Delayed Rewards (RUDDER) [43], which is a novel RL approach for delayed rewards in finite MDPs. RUDDER’s objective is to neutralize expected future rewards, thereby simplifying Q-value estimations to the average of immediate rewards. Despite RUDDER’s expedited processing in scenarios with delayed rewards compared to traditional RL methods, it fails to incorporate prior knowledge—an area where LMGT particularly excels. Consequently, we anticipate LMGT to facilitate the expedited development of effective behavioral strategies by agents. The second segment of the experiment evaluates our framework’s versatility by applying it across diverse RL algorithms and environments to ascertain its efficacy. Herein, we examined the enhancement in performance attributable to our framework across various RL algorithms, compared to some classic baselines. This phase further includes an assessment of the impact of different prompting techniques on our framework’s performance and an exploratory evaluation of the reasoning capabilities of LLMs within our framework. To ensure that the evaluation conclusions of our framework extend beyond synthetic settings, the final section investigates its application to and performance in complex robotic environments. We conducted experiments using the Housekeep environment [24]. Housekeep is a simulated environment for embodied agents, where robots must perform housekeeping tasks by correctly placing out-of-place objects into appropriate containers. To ensure that the evaluation conclusions of our framework extend beyond synthetic settings, the final section investigates its practical applications and improvements. Specifically, we explore its integration with Google’s SlateQ [25], a sophisticated recommendation algorithm that employs slate decomposition. This approach effectively manages the complexity of recommending multiple items simultaneously, addressing the challenge of large action spaces found in previous RL recommendation algorithms. This implementation was tested within a simulated environment on RecSim [44], a versatile platform for developing simulation environments for recommender systems (RSs), facilitating sequential user interactions.
4.1 Experimental Settings
@@
4.2 Comparison Experiments with Traditional Exploration-Exploitation Trade-off Methods
Reinforcement learning (RL) presents two fundamental challenges: the credit assignment problem for delayed rewards and the efficiency of exploration in complex environments. The former concerns the attribution of sparse, delayed rewards to preceding actions in a decision sequence, while the latter addresses the optimal balance between exploration and exploitation in unknown environments. Addressing these challenges has led researchers to develop various approaches, notably reward redistribution methods such as RUDDER and intrinsic motivation techniques such as NGU (Never Give Up). In this section, we empirically evaluate LMGT’s efficacy in addressing these core challenges through comparative experiments with these representative algorithms.
4.2.1 Comparison with RUDDER’s Reward Redistribution Method
RUDDER (Return Decomposition for Delayed Rewards) represents a specialized algorithm designed to address delayed reward problems. It employs a reward redistribution function that decomposes cumulative delayed rewards and attributes them to key preceding decision points. This approach is particularly effective for scenarios with challenging credit assignment, exemplified by the pocket watch repair task utilized in our experiments. In this task, agents must determine whether to repair watches of specific brands by conducting cost-benefit analyses between known selling prices and uncertain repair costs. The task embodies a classic delayed reward problem, as the profitability of repair decisions becomes apparent only after all associated costs are finalized.
Method Episodes Time (sec) TD 71,823 427 MC 221,770 530 RUDDER 2,029 171 LMGT+TD (ours) 417 114 Lower values indicate better performance in both metrics.
RUDDER was selected as a baseline for comparison because it represents a state-of-the-art method for addressing delayed reward problems and closely aligns with our research objectives. To ensure methodological consistency, LMGT was implemented with temporal difference (TD) learning, matching RUDDER’s foundational approach. Experiments were conducted using multiple runs with random seeds 42, 43, 44, 45, 46, and results were averaged. Performance was evaluated using two metrics: the number of training episodes required to achieve a profitable decision rate exceeding 90% (reported as “Episode”) and the total computational time needed for training (reported as “Time”), inclusive of LLM inference time.
Environment Algorithm Early Training Late Training 0 0.5 1.0 1.5 2.0 2.5 3.0 3.5 Pitfall R2D2 0.0 103.5 95.3 325.6 2786.3 3586.7 3548.6 3613.5 NGU+R2D2 0.0 786.9 3233.3 4020.2 6400.0 5494.6 5120.0 5973.4 LMGT+R2D2 (ours) 0.0 1535.5 5486.3 6348.2 6125.6 6354.3 6535.2 6503.5 Montezuma’s Revenge R2D2 0.0 1235.6 2567.2 2465.8 2500.7 2754.6 2703.5 2687.2 NGU+R2D2 0.0 4818.5 6369.2 7165.7 8236.5 9810.0 9966.7 9049.4 LMGT+R2D2 (ours) 0.0 5635.4 8957.3 11257.3 10985.7 12093.9 12463.6 12365.5 LMGT is our proposed method. Best results are highlighted in bold.
As illustrated in Table 1, experimental results demonstrate that LMGT substantially outperforms RUDDER on both metrics: LMGT requires only 417 training episodes and 114 seconds to develop a qualified strategy, whereas RUDDER necessitates 2029 episodes and 171 seconds. Notably, LMGT’s proportional advantage in training episodes (approximately 79.4% reduction) significantly exceeds its advantage in computational time (approximately 33.3% reduction). This disparity suggests that the prior knowledge embedded in the language model effectively accelerates the value learning process, resulting in more efficient credit assignment.
4.2.2 Comparison with NGU’s Exploration-Driven Method
With respect to the second critical challenge in reinforcement learning—exploration efficiency—NGU (Never Give Up) represents an advanced exploration methodology based on intrinsic motivation principles. NGU integrates novelty-driven intrinsic rewards with environmental extrinsic rewards, thereby encouraging continuous exploration of unknown state spaces while maintaining focus on high-value regions. While this approach demonstrates excellent performance in environments with sparse rewards, it nevertheless requires extensive environmental interactions to develop effective exploration strategies.
To assess LMGT’s exploration efficiency advantages, we selected two Atari games characterized by highly sparse rewards: Pitfall and Montezuma’s Revenge. These environments have been established as particularly challenging benchmarks for exploration algorithms due to their complex state spaces and extremely sparse reward mechanisms. For these experiments, R2D2 (Recurrent Replay Distributed DQN) served as the base algorithm, combined separately with NGU and LMGT. Performance was compared across various training frame quantities.
Experimental results (as shown in Figure 3 and Table 2) reveal that LMGT+R2D2 progressively achieves superior performance in both environments as training progresses. After 3.5×101̂0 frames of training, LMGT+R2D2 attains an average reward of 6503.5 in the Pitfall environment, surpassing NGU+R2D2 (5973.4) by approximately 8.9% and exceeding the baseline R2D2 implementation (3613.5) by approximately 80.0%. In the more challenging Montezuma’s Revenge environment, the performance differential increases markedly: LMGT+R2D2 achieves a reward of 12365.5, exceeding NGU+R2D2 (9049.4) by approximately 36.6% and outperforming baseline R2D2 (2687.2) by a factor of 4.6.
Perhaps more significant are the performance disparities observed during early training phases. With merely 0.5×101̂0 training frames, LMGT+R2D2 achieves a reward of 1535.5 in Pitfall, nearly double that of NGU+R2D2 (786.9). Similarly, in Montezuma’s Revenge, LMGT+R2D2 (5635.4) substantially outperforms NGU+R2D2 (4818.5). These early-stage advantages indicate that LMGT, leveraging prior knowledge embedded in language models, can more rapidly identify high-value regions, thereby significantly enhancing exploration efficiency.
4.2.3 Comprehensive Analysis and Discussion
Our comparative experiments with RUDDER and NGU demonstrate LMGT’s substantial advantages in addressing both core reinforcement learning challenges. These advantages derive from LMGT’s distinctive operational mechanism: language models not only provide structural environmental knowledge but also guide exploration through intelligent reward shifting, simultaneously optimizing both credit assignment processes and exploration strategies.
For delayed reward problems, LMGT utilizes language model-based action value assessment to provide more direct credit assignment mechanisms, enabling agents to rapidly identify critical decision points and substantially reducing required training episodes. For exploration challenges, LMGT employs language model evaluations of state-action pairs to assign positive reward shifts to potentially valuable trajectories, thus avoiding inefficient exploration and achieving more optimal exploration-exploitation balance in complex environments.
It is particularly noteworthy that LMGT’s advantages become increasingly pronounced as task complexity increases. This scalability arises because language models based on Transformer architectures maintain relatively constant inference speeds regardless of task complexity, whereas traditional methods typically require exponentially increasing exploration time as environments become more complex. This characteristic renders LMGT especially suitable for high-dimensional environments and long-horizon decision problems.
In conclusion, by integrating the cognitive capabilities of large language models with reinforcement learning frameworks, LMGT provides an innovative approach that simultaneously addresses credit assignment and exploration efficiency challenges. This integration opens new avenues for reinforcement learning applications in resource-constrained scenarios, potentially transforming how complex sequential decision problems are approached.